# Insight Over Sight? Exploring the Vision-Knowledge Conflicts in Multimodal LLMs

### v1

@@ Abstract
This paper explores the problem of commonsense-level vision-knowledge conflict in Multimodal Large Language Models (MLLMs), where visual information contradicts model’s internal commonsense knowledge (see Figure 1). To study this issue, we introduce an automated pipeline, augmented with human-in-the-loop quality control, to establish a benchmark aimed at simulating and assessing the conflicts in MLLMs. Utilizing this pipeline, we have crafted a diagnostic benchmark comprising 374 original images and 1,122 high-quality question-answer (QA) pairs. This benchmark covers two types of conflict targets and three question difficulty levels, providing a thorough assessment tool. Through this benchmark, we evaluate the conflict-resolution capabilities of nine representative MLLMs across various model families and find a noticeable over-reliance on textual queries. Drawing on these findings, we propose a novel prompting strategy, “Focus-on-Vision” (FoV), which markedly enhances MLLMs’ ability to favor visual data over conflicting textual knowledge. Our detailed analysis and the newly proposed strategy significantly advance the understanding and mitigating of vision-knowledge conflicts in MLLMs. The data and code are made publicly available. [^1] https://github.com/xyliu-cs/ConflictVIS
@@ 1 Introduction
Large Language Models (LLMs) (Brown et al., 2020; OpenAI, 2022; Touvron et al., 2023; OpenAI, 2023a; Meta AI, 2024) have reshaped the landscape of deep learning for their comprehensive capabilities in language understanding, reasoning, and generation (Wei et al., 2022a; b; Chen et al., 2023). This evolution has paved the way for the emergence of Multimodal Large Language Models (MLLMs) (Li et al., 2022a; 2023b; Lyu et al., 2023; Dai et al., 2023; Liu et al., 2024d; Zhu et al., 2023; Bai et al., 2023; Liu et al., 2024b; Wang et al., 2024; OpenAI, 2023b; Liu et al., 2024c; Tong et al., 2024; Gemini Team, 2024; OpenAI, 2024b), which integrate a vision model with the LLM to process and respond to visual information. MLLMs like GPT-4o (OpenAI, 2024b) and LLaVA-NeXT (Liu et al., 2024c) have exhibited remarkable proficiency across various vision-language tasks such as image captioning (Chen et al., 2015), visual question answering (VQA) (Antol et al., 2015), and visual reasoning (Johnson et al., 2017; Yue et al., 2024).
However, the persistence of knowledge conflicts in LLMs remains a significant challenge for MLLMs, often cited as a primary cause of hallucinations in these systems (Liu et al., 2023a; Guan et al., 2024). Introducing visual information into MLLMs generates a novel form of discrepancy termed “vision-knowledge conflict”, where the visual data contradicts the model’s pre-existing parametric knowledge. Prior research has mainly focused on assessing and comprehending these conflicts through the lens of factual knowledge by generating counterfactual images via image editing techniques (Guan et al., 2024). Nonetheless, the aspect of vision-knowledge conflict related to commonsense knowledge, [^2] To illustrate, a factual knowledge example is “The Eiffel Tower is 330 meters tall”, while a commonsense knowledge example is “Babies cry when they are hungry”. which proves more challenging for LLMs and MLLMs due to its implicit and nuanced nature (Li et al., 2022b), remains under-investigated.
@@
To address this gap, we propose an automated pipeline with human-in-the-loop to develop a benchmark for simulating and analyzing commonsense-level vision-knowledge conflicts in MLLMs. Illustrated in Figure 2, our pipeline consists of four key modules: knowledge component extraction, counter-commonsense query construction, image generation, and question-and-answer (QA) generation. This framework streamlines the generation of counter-commonsense inputs from scratch and is modularly designed to facilitate the addition of further conflict categories and QA formats in the future. We demonstrate the application of our framework by developing the ConflictVis benchmark built on the Open Mind Common Sense dataset (Singh et al., 2002). The benchmark focuses on the critical aspects of commonsense knowledge in the visual scene: Subject, Action, and Place. It consists of 374 original images with 1,122 high-quality QA pairs spanning two conflict targets and three question difficulty levels, all verified by human annotators.
@@
Using the crafted benchmark, we evaluate nine representative MLLMs from five model families, providing insights into model behaviors, the causes of conflicts, and effective methods designs to mitigate the negative effects of conflicts. Notably, when facing knowledge conflicts, MLLMs tend to over-rely on their parametric knowledge for simpler questions. For instance, the top commercial model, Claude-3.5-Sonnet, exhibits a memorization ratio (MR) on parametric knowledge of 43.6% on Yes-No questions, substantially higher than results on more complex subjective questions; the largest open-source model evaluated, LLaVA-NeXT-34B, shows a MR of 26.7% on Yes-No questions, three times more than that on subjective questions. Our detailed analysis of the failure cases reveals that MLLMs generally focus more on the textual query than on the visual context. Drawing on these observations, we revisit several existing training-free methods (Leng et al., 2024; Liu et al., 2024f; Wei et al., 2022b) to enhance the impact of visual context in answer generation. Interestingly, although Chain-of-Thought prompting (Wei et al., 2022b) improves model’s reasoning abilities, it directs MLLMs to over-utilize the parametric knowledge in the reasoning process, leading to inaccurate conclusions or refusals. From this understanding, we propose Focus-on-Vision (FoV) prompting to instruct MLLMs to prioritize visual information, which markedly improves the performance.
Our main contributions are summarized as follows:
@@
  • We introduce an innovative framework to automatically construct counter-commonsense benchmarks from the ground up. This platform allows the flexible definition of conflict categories and QA formats, facilitating the large-scale creation of conflict samples with minimal human effort.
@@
  • We present ConflictVis , a pioneering diagnostic benchmark specifically designed to evaluate vision-knowledge conflicts at the commonsense level in MLLMs. The benchmark is meticulously validated by human experts to guarantee the quality of data.
  • We perform extensive experiments on nine representative MLLMs, uncovering critical insights into model performance and the underlying mechanisms of the conflict. Leveraging these insights, we propose a novel Focus-on-Vision (FoV) prompting as an effective solution to significantly minimize vision-knowledge conflicts in MLLMs.
@@ 2 ConflictVis Benchmark
Figure 2: An automated pipeline with human quality control to construct images and question-answer pairs that conflict with commonsense knowledge.
@@
This section outlines the pipeline and the data utilized to construct the ConflictVis benchmark.
2.1 Automated Pipeline
@@
Commonsense knowledge refers to the information generally accepted by the majority of people about everyday life, encompassing practical principles on how the world functions (Singh et al., 2002). Inspired by the widely-used image captioning dataset (Chen et al., 2015), we explore a specific type of commonsense knowledge encapsulated as a triplet $<s,a,p>$, where $s$ is the Subject, $a$ is the Action, and $p$ is the Place of the action or subject. For example, the statement “a baby (Subject) drinking water (Action) on the bed (Place)” illustrates this format. In this structure, the Subject outlines the main object’s appearance, the Action describes the primary activity and its interactions with other relevant objects, and the Place highlights the background objects and setting. This three-part framework efficiently captures the vital details that are accurately reflected in an image.
We then describe the strategies for creating images and corresponding QA pairs that contradict the commonsense knowledge based on an input corpus. In general, we construct triplets containing low co-occurring concepts to serve as the counter-commonsense queries for image generation and fill pre-defined templates to generate the QA pairs. The pipeline contains the following four steps:
1. Extract Knowledge Components from Input Corpus.
The process commences with the extraction of knowledge components from the corpus, which contains human commonsense knowledge in natural language. Drawing inspiration from Li et al. (2021) on the creation of a compositional generalization test set, we identify frequently occurring components in the corpus and enumerate all possible combinations to generate knowledge triplets. To this end, we employ a transformer model pipeline to extensively annotate the syntactic labels, including dependency (DEP), Part-of-Speech (POS), and Named Entity (NE). Table 3 in Appendix lists the linguistic rules to identify each component:
  • Action Phrases: The process starts by identifying the verb token (“VERB”) and then examines its dependents for direct objects marked as “dobj”. Upon finding such objects, our framework includes any determiners or compound modifiers present to form the complete action phrase. In addition, verbs are converted to the "VBG" form to align with the common practice in image captions.
  • Place Phrases: Distinguishing place phrases from other prepositional phrases relies on more than just syntactic labels. Accordingly, the framework employs semantic role labeling techniques by utilizing a BERT-based model (Gardner et al., 2017) to annotate each word’s semantic role. It then extracts the phrase with a location tag (“ARGM-LOC") that contains 3 to 4 words.
To develop a high-quality benchmark, our framework filters the top subject ($N_{S}$), action ($N_{A}$), and place ($N_{P}$) phrases, respectively, based on the occurrence. It also removes named entities from the subject and place phrases and merges similar phrases (e.g., “a doctor” and “the doctor”) by keeping the most frequently occurring variant.
2. Construct Counter-commonsense Query.
We aim to construct scenes with a single anomalous concept (i.e., target) that seldom co-occurs with the other concepts (i.e., the context pair). For instance, the Action component “fixing a computer” and the context pair (“the baby”, “on the bed”) results in the triplet “the baby on the bed fixing a computer”, which contradicts commonsense knowledge, as it is highly improbable for a baby to fix a computer in reality. Specifically,
  • The context pair should exhibit a high co-occurrence level, acting as a common background.
  • The target should display a low co-occurrence with the given context, representing the anomaly.
To find the high co-occurrence context pairs, our framework first groups context pairs by their target category, i.e., ContextAction consists of (Subject, Place) pairs and ContextPlace comprises (Subject, Action) pairs. These groups help to develop the focus of the questions in the next section. We omit ContextSubject to avoid the ambiguity and potential ethical issues for judging a subject’s identity.
We demonstrate how to construct queries with counter-commonsense Actions, and a similar approach can be applied to Places. For each subject, we pair it with all the Place components to establish context candidates and select the contexts based on the co-occurrence level. We use Normalized Pointwise Mutual Information (NPMI) to evaluate the co-occurrence of the context pairs ${\bf C}=(C_{X},C_{Y})$:
$\text{PMI}(C_{X};C_{Y})\equiv\log_{2}\frac{P(C_{X},C_{Y})}{P(C_{X})P(C_{Y})},\qquad\text{NPMI}(C_{X};C_{Y})\equiv\frac{\text{PMI}(C_{X};C_{Y})}{-\log_{2}P(C_{X},C_{Y})}.$ (1)
A higher NPMI score indicates a more frequent co-occurrence of the two context components. For each subject, we select Top-$K$ places with highest NPMI scores to form context pairs.
For each selected context pair, we construct counter-commonsense triplets by selecting a target Action component that is not compatible with the context pair. Concretely, we also use the NPMI metric (i.e. $\text{NPMI}(T;{\bf C})$) to evaluate the co-occurrence of the target $T$ and context pair ${\bf C}$. For each context pair, we select Top-$M$ actions with lowest NPMI scores to form counter-commonsense queries.
To better match real-world knowledge, we leverage an LLM trained on large-scale web data to calculate the probability $P(\cdot)$. For example, to calculate the joint probabilities for multiple co-occurring components,our framework forms a concatenated expression of the concept phrases (e.g., “a baby on the bed” or “a baby on the bed fixing a computer”) and inputs it into a pre-trained LLM to obtain the generative probability. Utilizing an LLM here offers an advantage over frequency counting for addressing the issue of zero frequencies for rare concept compositions, which often arises with limited dataset size. Since large language models are trained on large-scale data to learn the statistical tendencies of human language, their generative probability provides an approximation of the likelihood of certain phrase appearing in reality. For the probability of each individual component, we use its relative frequency within its category (Lin, 1999). Following generation, a manual review is conducted here to ensure the data quality (See Appendix B for further details).
3. Generate Image.

### v2

# Insight Over Sight: Exploring the Vision-Knowledge Conflicts in Multimodal LLMs

@@ Abstract
This paper explores the problem of commonsense level vision-knowledge conflict in Multimodal Large Language Models (MLLMs), where visual information contradicts model’s internal commonsense knowledge. To study this issue, we introduce an automated framework, augmented with human-in-the-loop quality control, to generate inputs designed to simulate and evaluate these conflicts in MLLMs. Using this framework, we have crafted a diagnostic benchmark consisting of 374 original images and 1,122 high-quality question-answer (QA) pairs. The benchmark covers two aspects of conflict and three question types, providing a thorough assessment tool. We apply this benchmark to assess the conflict-resolution capabilities of nine representative MLLMs from various model families. Our results indicate an evident over-reliance on parametric knowledge for approximately 20% of all queries, especially among Yes-No and action-related problems. Based on these findings, we evaluate the effectiveness of existing approaches to mitigating the conflicts and compare them to our “Focus-on-Vision” prompting strategy. Despite some improvement, the vision-knowledge conflict remains unresolved and can be further scaled through our data construction framework. Our proposed framework, benchmark, and analysis contribute to the understanding and mitigation of vision-knowledge conflicts in MLLMs.
Insight Over Sight: Exploring the Vision-Knowledge Conflicts
in Multimodal LLMs
Xiaoyuan Liu1,2thanks:   Work was done when Xiaoyuan, Wenxuan, Youliang, and Jen-tse were interning at Tencent.    Wenxuan Wang3    Youliang Yuan1,2    Jen-tse Huang4    Qiuzhi Liu2    Pinjia He1thanks:   Pinjia He is the corresponding author.    Zhaopeng Tu2 1School of Data Science, The Chinese University of Hong Kong, Shenzhen 2Tencent    3Renmin University of China    4Johns Hopkins University 1xiaoyuanliu@link.cuhk.edu.cn     1hepinjia@cuhk.edu.cn     2zptu@tencent.com
1 Introduction
@@
Large Language Models (LLMs) (Brown et al., 2020; OpenAI, 2022; Touvron et al., 2023; OpenAI, 2023a; Meta AI, 2024) have reshaped the landscape of deep learning for their comprehensive capabilities in language understanding, reasoning, and generation (Wei et al., 2022a, b; Chen et al., 2023). This evolution has paved the way for the emergence of Multimodal Large Language Models (MLLMs) (Li et al., 2022, 2023b; Lyu et al., 2023; Dai et al., 2023; Liu et al., 2024d; Zhu et al., 2023; Bai et al., 2023; Liu et al., 2024b; OpenAI, 2023b; Liu et al., 2024c; Tong et al., 2024; Gemini Team, 2024; OpenAI, 2024b; Meta AI, 2025), which integrate a vision model with the LLM to process visual information. Modern MLLMs like GPT-4o (OpenAI, 2024b) and LLaVA-NeXT (Liu et al., 2024c) have exhibited remarkable proficiency across various vision-language tasks such as image captioning (Chen et al., 2015), visual question answering (VQA) (Antol et al., 2015), and visual reasoning (Johnson et al., 2017; Yue et al., 2024).
However, the persistence of knowledge conflicts in LLMs remains a significant challenge for MLLMs. Introducing visual information into MLLMs generates a novel form of discrepancy, a phenomenon we term “vision-knowledge conflict”, in which the visual data contradicts the model’s pre-existing parametric knowledge. Previous studies have demonstrated that LLMs can exhibit complex behavior when confronted with knowledge conflicts, oscillating between rigid adherence to their parametric knowledge and excessive sensitivity to contextual cues (Xie et al., 2024). This behavior brings substantial risks, especially when external information is critical for decision-making. Given the growing need for trustworthiness, real-time accuracy, and robustness in MLLM systems, it is imperative to further investigate and resolve these vision-knowledge conflicts, which are believed to be a primary cause of hallucinations (Liu et al., 2023a; Guan et al., 2024). While recent research has simulated these conflicts by manually crafting counterfactual images (Guan et al., 2024), there remains significant room for improvement in conflict categorization, question diversity, and image naturalness. Additionally, manually crafted benchmarks are limited in sample size and scalability, underscoring the need for a more automated pipeline to enable broader analysis.
@@
To this end, we propose an automated framework with human-in-the-loop to develop a benchmark for simulating and analyzing vision-knowledge conflicts in MLLMs at the commonsense level  [^1] For example, a commonsense knowledge statement is “babies cry when they are hungry”, as opposed to a factual knowledge statement like “the Eiffel Tower is 330 meters tall.”. Illustrated in Figure 2, our framework consists of 3 key modules: (1) knowledge component extraction, (2) counter-commonsense query construction, and (3) image and question-answer (QA) pair synthesis. This framework streamlines the generation of counter-commonsense inputs from scratch and is modularly designed to facilitate the addition of new conflict categories and QA formats in the future. We demonstrate the application of our framework by developing the ConflictVis benchmark, focusing on the aspects of Subject, Action, and Place. The benchmark consists of 374 original images with 1,122 high-quality QA pairs spanning two conflict targets and three question types, all manually verified.
@@
Using the crafted benchmark, we evaluate nine representative MLLMs from five model families, providing insights into model behaviors, the causes of conflicts, and effective approaches to mitigate the negative effects of conflicts. Notably, when facing knowledge conflicts, MLLMs tend to over-rely on their parametric knowledge for the answer, especially among Yes-No questions and action-related problems. For instance, the top commercial model, Claude-3.5-Sonnet, exhibits a memorization ratio on parametric knowledge of 43.6% on Yes-No questions, substantially higher than results on more complex Open-Ended questions. Regarding conflict type, the average memorization ratio for counter-commonsense action problems is 23.8%, 10.4 percentage points higher than that for the place problems. Our detailed analysis of the failure cases reveals that MLLMs generally underutilize visual information, relying on parametric knowledge to infer the answer based on textual clues. Drawing on these observations, we assess several existing improvement methods to enhance the impact of visual context in answer generation. Interestingly, although Chain-of-Thought prompting improves the reasoning abilities, it guides MLLMs to utilize parametric knowledge more during rationalization, often resulting in contradictory conclusions or refusals. In response, we propose “Focus-on-Vision” (FoV) prompting to directly instruct MLLMs to prioritize visual information, which markedly improves the model’s performance. Despite advancements by various mitigation approaches, the vision-knowledge conflict remains a persistent challenge.
Our main contributions are summarized below:
@@
  • We introduce an innovative framework to automatically construct counter-commonsense benchmarks from the ground up. This framework allows the flexible definition of conflict categories and QA formats, facilitating the large-scale creation of conflict samples with minimal human effort.
@@
  • We present ConflictVis, a pioneering diagnostic benchmark specifically designed to evaluate the commonsense level vision-knowledge conflicts in MLLMs. The benchmark is meticulously validated by human experts to guarantee the quality of data.
  • We benchmark nine representative MLLMs and evaluate the effectiveness of several improvement methods in addressing conflicts. Through this analysis, we demonstrate the significance of the vision-knowledge conflict.
2 Related Work
Knowledge Conflicts.
Knowledge conflicts in LLMs can be divided into three categories (Xu et al., 2024c): within the retrieved context (Chen et al., 2022), within model’s parametric knowledge (Huang et al., 2023), and between the context and the model’s parametric knowledge (Xie et al., 2024; Wu et al., 2024a; Su et al., 2024). These conflicts can lead to incorrect or inconsistent responses, undermining the model’s trustworthiness (Xu et al., 2024c; Xie et al., 2024). In MLLMs, these conflict types expand to multimodal inputs, where conflicts can occur between the input image and the text instruction (Liu et al., 2024e; Han et al., 2024), or when the image contains counterfactual information that contradicts the model’s parametric knowledge (Liu et al., 2024e; Guan et al., 2024).
To evaluate the conflicts between visual information and the parametric knowledge in MLLMs, HallusionBench (Guan et al., 2024) consists of manually edited informational graphics, forming (normal, counterfactual) image pairs used to assess model consistency through Yes-No questions. AutoHallusion (Wu et al., 2024b) introduces an automated approach to generate counterfactual scenarios by altering correlated objects in images, utilizing Yes-No questions to probe object existence and spatial relationships. In the context of commonsense knowledge, PhD (Liu et al., 2024e) generates counter-commonsense images through manual collection and synthesis, employing short-answer questions for assessment. In contrast, our proposed benchmark features an automated framework and a broad array of question types and conflict targets, enabling a more scalable and comprehensive evaluation of MLLMs.
Hallucination in MLLMs.
Hallucination in MLLMs refers to the situation where the model generates descriptions that conflict with the given visual context. MLLMs’ hallucinations are typically categorized based on the type of incorrect information, such as non-existent objects, incorrect object attributes, and inaccurate object relations (Liu et al., 2024a). Relevant research has mainly focused on two key areas: developing benchmarks and metrics to detect hallucinations (Rohrbach et al., 2018; Li et al., 2023c; Liu et al., 2023a; Wang et al., 2023) and proposing strategies to mitigate them (Liu et al., 2023a; Leng et al., 2024; Huang et al., 2024; Liu et al., 2024f).
In the context of knowledge conflicts, hallucination occurs when the model gives precedence to its intrinsic knowledge rather than the visual context (Liu et al., 2023a; Guan et al., 2024; Liu et al., 2024e, a). Our study explores this issue by simulating knowledge conflicts and evaluating the effectiveness of various hallucination-mitigation approaches in resolving the conflicts.
Benchmarks for MLLMs.
Traditional vision-language benchmarks are designed to assess independent skills, including image captioning (Chen et al., 2015), visual grounding (Rohrbach et al., 2016), and visual question answering (Antol et al., 2015; Hudson and Manning, 2019). However, with the emergence of MLLMs, there is a growing need for more comprehensive and tailored benchmarks. The strong zero-shot abilities and advanced language generation skills exhibited by MLLMs make traditional benchmarks insufficient, as they may not account for the diversity of responses or the full range of MLLM capabilities. To address these limitations, researchers have developed more complex benchmarks to evaluate MLLMs across a wider range of tasks (Yue et al., 2024; Fu et al., 2023; Li et al., 2023a; Liu et al., 2023c). Meanwhile, diagnostic benchmarks have also been built to evaluate specific challenges or traits in MLLMs, such as hallucination (Li et al., 2023c; Liu et al., 2023a), social bias (Howard et al., 2024), and model safety (Liu et al., 2023b). ConflictVis is a pioneering analytical benchmark that presents conflicting visual contexts to challenge the model’s commonsense knowledge, enabling deeper investigations into model performance in the presence of vision-knowledge conflicts.
@@ 3 ConflictVis Benchmark
Figure 2: An automated framework with human quality control to construct images and question-answer pairs that conflict with commonsense knowledge.
@@
This section outlines the framework and the data utilized to construct the ConflictVis benchmark.
3.1 Automated Framework
@@
Commonsense knowledge refers to the information generally accepted by the majority of people about everyday life, encompassing practical principles on how the world functions (Singh et al., 2002). Based on the widely-used image captioning dataset (Chen et al., 2015), we explore a specific type of commonsense knowledge encapsulated as a triplet $\langle\bm{s},\bm{a},\bm{p}\rangle$, where $s$ is the Subject, $a$ is the Action, and $p$ is the Place of the action or subject. For example, the statement “the waitress (Subject) washing dishes (Action) in the kitchen (Place)” illustrates this format. In this structure, the Subject outlines the main object’s appearance, the Action describes the primary activity and its interactions with other relevant objects, and the Place highlights the background objects and setting. This three-part framework efficiently captures the vital details that are accurately reflected in an image.
We next describe our approach for generating images and corresponding QA pairs that challenge the commonsense knowledge in MLLMs. In general, we construct triplets of low co-occurring concepts to serve as counter-commonsense queries for the multimodal input generation. Our framework includes three stages, as depicted in Figure 2. We detail each stage below.
Extract Knowledge Components
Drawing inspiration from Li et al. (2021) on the creation of a compositional generalization test set, we identify Subject, Action, and Place phrases as the most frequent elements in the corpus and build compounds based on them. To this end, we first employ a transformer model pipeline Honnibal et al. (2020) to extensively annotate the syntactic labels, including dependency (DEP), Part-of-Speech (POS), and Named Entity (NE). Then, we extract Subject, Action, and Place phrases from the corpus based on the predefined linguistic rules (See Appendix A.1). To ensure the data quality, our framework selects the top $N$ phrases from each category based on the frequency. It further refines the data by removing named entities from Subject and Place phrases and consolidating similar expressions (e.g., “a doctor” and “the doctor”) by retaining the most frequently occurring variant. This process results in three curated phrase lists for subsequent processing.
Construct Counter-commonsense Query
We aim to construct scenes with one single anomalous component (i.e., target) that seldom co-occurs with the others (i.e., the context). This objective can be formalized into two key requirements: (1) The context components should exhibit a high co-occurrence level, acting as a common background, and (2) The target component should display a low co-occurrence with the given context, representing the anomaly. To identify strongly co-occurring context pairs, our framework first groups context components by the target category, e.g., ContextAction consists of (Subject, Place) pairs. These groupings help to develop the focus of the questions in the next section. We omit ContextSubject to prevent ambiguity and mitigate potential ethical concerns related to subject identity. Next, our framework enumerate all the context combinations within the each group and computes their Normalized Pointwise Mutual Information (NPMI) scores.
$\text{PMI}(C_{X};C_{Y})\equiv\log_{2}\frac{P(C_{X},C_{Y})}{P(C_{X})P(C_{Y})}$ (1)
$\text{NPMI}(C_{X};C_{Y})\equiv\frac{\text{PMI}(C_{X};C_{Y})}{-\log_{2}P(C_{X},C_{Y})}$ (2)
A higher NPMI score indicates a stronger co-occurrence between the two components. For each context group, the framework selects Top-$K$ context pairs with the highest NPMI scores to form a candidate pool. For each context pair in the pool, the framework generates counter-commonsense triplets by selecting a target component that is unusual given the context. Concretely, the framework evaluates the co-occurrence between each target $T$ and the context pair ${\bf C}$ using the NPMI score (i.e. $\text{NPMI}(T;{\bf C})$). For each context pair, the Top-$M$ targets with the lowest NPMI scores are retained to construct counter-commonsense queries. To better align the model knowledge, the framework queries an LLM trained on large-scale web data to estimate the probability $P(\cdot)$, and a manual review is conducted after the query generation to ensure quality (See Appendix A.2 for more details).
Generate Multimodal Inputs
Building on the constructed queries, our framework generates three types of questions, Yes-No, Multiple-Choice, and Open-Ended, along with their corresponding answers. To achieve this, the framework employs predefined question templates and fills in the relevant components accordingly. Example question-answer pairs are provided in Table 1. To generate corresponding images, our framework concatenates the triplet components into a caption-like expression and employs a prompt template to query a text-to-image model, as illustrated in Figure 2. Following image generation, human annotators perform quality control by filtering out low-quality images, such as those that appear distorted or misaligned with the input prompt. Detailed image filtering guidelines, along with illustrative examples, are provided in the Appendix B.3.
Question Content Answer
Yes-No Is the waitress in the kitchen signing a bill? Yes
Multiple-Choice Question: What is the waitress doing in the kitchen?
Options: (A) washing dishes. (B) riding a bike. (C) starting a fire. (D) signing a bill.
D
Open-Ended What is the waitress doing in the kitchen? signing a bill
Table 1: Example question-answer pairs generated by our framework.
3.2 ConflictVis Benchmark Construction